OpenStudy (anonymous):

Hello, i need help to solve by Bernouilli the ode's xy'-4x^2 y+2ylny i'll thanks forever

OpenStudy (anonymous):

What is this equal to? 0?

OpenStudy (anonymous):

divide the whole equation by y then put lny=t (a variable) differentiating this substitution with respect to x results in {1/y}*{dy/dx}=dt/dx putting this value in your question will reduce it to LINEAR DIFFERENTIAL EQUATION which can be easily be solved